# When Attention Sink Emerges in Language Models: An Empirical View

### v1

@@ Abstract
Language Models (LMs) assign significant attention to the first token, even if it is not semantically important, which is known as attention sink. This phenomenon has been widely adopted in applications such as streaming/long context generation, KV cache optimization, inference acceleration, model quantization, and others. Despite its widespread use, a deep understanding of attention sink in LMs is still lacking. In this work, we first demonstrate that attention sinks exist universally in LMs with various inputs, even in small models. Furthermore, attention sink is observed to emerge during the LM pre-training, motivating us to investigate how optimization, data distribution, loss function, and model architecture in LM pre-training influence its emergence. We highlight that attention sink emerges after effective optimization on sufficient training data. The sink position is highly correlated with the loss function and data distribution. Most importantly, we find that attention sink acts more like key biases, storing extra attention scores, which could be non-informative and not contribute to the value computation. We also observe that this phenomenon (at least partially) stems from tokens’ inner dependence on attention scores as a result of softmax normalization. After relaxing such dependence by replacing softmax attention with other attention operations, such as sigmoid attention without normalization, attention sinks do not emerge in LMs up to 1B parameters. The code is available at https://github.com/sail-sg/Attention-Sink.
@@ 1 Introduction
Xiao et al. (2023) showed that Large Language models (LLMs) allocate significant attention to the initial tokens, irrespective of their semantic relevance. This interesting phenomenon is termed as attention sink and has widespread applications, including streaming/long context generation (Xiao et al., 2023; Han et al., 2024; Yang et al., 2024), KV cache optimization (Ge et al., 2023; Wan et al., 2024; Wu & Tu, 2024), efficient inference (Zhang et al., 2024; Chen et al., 2024), model quantization (Liu et al., 2024b; Huang et al., 2024), and others.
A seminal of works attempted to understand attention sink. Among them, Cancedda (2024) clarified that attention sink primarily appears only on the first token. They found that early FFNs in LLaMA2 blast off the large norm of hidden states of the first token, thus leading to the attention sink in later layers. This is referred to as massive activations (very few activations exhibit extremely large values compared to others) in Sun et al. (2024). Besides, Sun et al. (2024); Yu et al. (2024) observed that attention sink may also appear in several word tokens carrying limited semantic information and having no fixed position. Despite the above research efforts, a deep understanding of attention sink is still absent. Therefore, we conduct a comprehensive study to investigate when attention sink emerges.
Based on open-sourced LLMs, we show that the first token acts as biases: the angles between the first key and queries of other tokens are typically small, leading to attention sink. Then we find that attention sink universally exists in auto-regressive LMs across different inputs, even in the small models or with random token sequences. Additionally, attention sink is observed to emerge during the LM pre-training before continual instruction tuning (Ouyang et al., 2022). This motivates us to focus on the LM pre-training, whose objective can be formulated as:
@@ 2 Preliminaries on LMs and attention sink
$\bm{A}^{l\textrm{,}h}=\textrm{Softmax}\big{(}{\bm{Q}^{l\textrm{,}h}{\bm{K}^{l\textrm{,}h}}^{T}}/{\sqrt{d_{h}}}+\bm{M}\big{)}\textrm{,}\,\,\,\,\bm{O}^{l}=\textrm{Concat}_{h=1}^{H}\left(\bm{A}^{l\textrm{,}h}\bm{V}^{l\textrm{,}h}\right)\bm{W}_{O}^{l}\textrm{,}$ (4)
@@
Positional embedding. NoPE (Kazemnejad et al., 2024) considered no explicit positional embedding (PE) in LMs, where $\bm{P}=\bm{0}$. When using absolute PE (Vaswani et al., 2017), $\bm{P}$ is a periodic function of token positions. Devlin et al. (2019); Brown et al. (2020) adopted a learnable PE, which means $\bm{P}$ is a learnable embedding of token positions. The dot product between each query and key meets $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{T}$ when using the above three PEs. While for relative PE (Raffel et al., 2020), ALiBi (Press et al., 2021), Rotary (Su et al., 2024), they have $\bm{P}=\bm{0}$. Instead, they modify the dot product $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle$. For relative PE and ALiBi, $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{T}+g(i-j)$, where $g(\cdot)$ is pre-defined function of the distance between two token positions. For Rotary, $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,j-i}\bm{k}_{j}^{T}$, where $\bm{R}_{\Theta\textrm{,}\,(\cdot)}$ is a pre-defined rotation matrix. We include detailed formulations of the above PEs in Appendix A.
@@
Attention sink. Xiao et al. (2023) revealed that LLMs allocate significant attention scores to specific token positions, e.g. the first token, resulting in “vertical” attention patterns. This holds even when the first token is not a BOS token (Xiao et al., 2023; Sun et al., 2024). To represent this, we have the attention scores $\bm{A}_{i\textrm{,}\,1}^{l\textrm{,}\,h}\gg\textrm{mean}(\bm{A}_{i\textrm{,}\,j\neq 1}^{l\textrm{,}\,h})$.
@@ 3 Properties of attention sink
Figure 2: In LLaMA3-8B Base, (Top) the first token has significantly larger $\ell_{2}$-norm of hidden states, but much smaller $\ell_{2}$-norm of keys and values than the mean of other tokens; (Bottom) cosine similarity instead of norm product contributes to attention sink. We delay more visualizations to Appendix B.3.
@@ 3.1 The first token acts as biases
Uniqueness of the first token. It is noted that the calculation of hidden states for the first token has no involvement of self-attention $\bm{h}_{1}^{l}=\textrm{FFN}(\textrm{LN}(\bm{o}_{1}^{l}+\bm{h}_{1}^{l-1}))+\bm{o}_{1}^{l}+\bm{h}_{1}^{l-1}$, where $\bm{o}_{1}^{l}=\textrm{LN}(\bm{h}_{1}^{l-1})\begin{bmatrix}\bm{W}^{l\textrm{,}1}&\bm{W}^{l\textrm{,}2}&\cdots&\bm{W}^{l\textrm{,}H}\\ \end{bmatrix}\bm{W}_{O}^{l}$. Therefore, $\bm{h}_{1}^{l}$, and corresponding queries/keys/values $\bm{k}_{1}^{l\textrm{,}h}=\textrm{LN}(\bm{h}_{1}^{l-1})\bm{W}_{K}^{l\textrm{,}h}$, $\bm{q}_{1}^{l\textrm{,}h}=\textrm{LN}(\bm{h}_{1}^{l-1})\bm{W}_{Q}^{l\textrm{,}h}$, $\bm{v}_{1}^{l\textrm{,}h}=\textrm{LN}(\bm{h}_{1}^{l-1})\bm{W}_{V}^{l\textrm{,}h}$ could be considered as the MLP output of input word embedding $\bm{x}_{1}\bm{W}_{E}$. Cancedda (2024); Sun et al. (2024) showed that $\bm{h}_{1}^{l}$ has large norm or massive activations, thus behaving as a bias to absorb the attention. Using LLaMA3-8B Base model (Dubey et al., 2024), we show that from certain transformer block, e.g., $l=2$, the $\ell_{2}$-norm of $\bm{h}_{1}^{l}$ is significantly larger than that of other tokens $\bm{h}_{t\neq 1}$ in Figure 2(Top). Despite the large $\ell_{2}$-norm of hidden states, we observe that the $\ell_{2}$-norm of keys and queries of the first token is significantly smaller than that of other tokens in the same figure. This motivates us to further explore the functionality of massive activations in hidden states.
@@
For the $t$-th query, we know that $\bm{q}_{t}^{l\textrm{,}\,h}{\bm{k}_{1}^{l\textrm{,}\,h}}^{T}$ typically has much larger values than $\bm{q}_{t}^{l\textrm{,}\,h}{\bm{k}_{j\neq 1}^{l\textrm{,}\,h}}^{T}$, as visualized in Figure 2(Bottom). We further show that due to the different manifold of $\bm{k}_{1}^{l\textrm{,}\,h}$, the angles between $\bm{k}_{1}^{l\textrm{,}\,h}$ and $\bm{q}_{t}^{l\textrm{,}\,h}$ play an important role. Considering $\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{j}^{l\textrm{,}h}}^{T}=\left\|\bm{q}_{t}^{l\textrm{,}h}\right\|\cdot\left\|\bm{k}_{j}^{l\textrm{,}h}\right\|\cdot\textrm{cos}(\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h})$, we visualize the cosine similarity between keys and values, and the product of $\ell_{2}$-norm between keys and values in Figure 2(Bottom). Although $\left\|\bm{q}_{t}^{l\textrm{,}h}\right\|\cdot\left\|\bm{k}_{1}^{l\textrm{,}h}\right\|$ is comparatively small, $\textrm{cos}(\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h})$ is significantly large, leading to attention sink. This explains why attention sink exists despite the small $\ell_{2}$-norm of keys of the first token. To conclude, the first token leverages its keys to act as biases, thus minimizing the angles between $\bm{k}_{1}^{l\textrm{,}\,h}$ and $\bm{q}_{t}^{l\textrm{,}\,h}$ and exhibiting attention sink.
@@ 3.2 Measuring attention sink
Threshold-based metrics. Xiao et al. (2023) showcased the appearance of attention sink by visualizing attention logits/scores in different heads/blocks. This leads to the intractability of measuring attention sink quantitatively due to the large number of attention heads and blocks. Therefore, we first explore the metrics to measure the attention sink. Within each head, we compute the importance scores for the $k$-th token $\alpha_{k}^{l\textrm{,}h}=\frac{1}{T-k+1}\sum_{i=k}^{T}\bm{A}_{i\textrm{,}\,k}^{l\textrm{,}\,h}$. We mainly focus on the first token $\alpha_{1}^{l\textrm{,}h}$. It is noted that $\frac{1}{T}\leq\alpha_{1}^{l\textrm{,}\,h}\leq 1$ since $\bm{A}_{1\textrm{,}\,1}^{l\textrm{,}\,h}=1$ and $0\leq\bm{A}_{i\neq 0\textrm{,}\,1}^{l\textrm{,}\,h}\leq 1$. Then we adopt a threshold-based metric, we consider a head has attention sink in the first token if $\alpha_{1}^{l\textrm{,}h}>\epsilon$. Considering that the whole model has $L$ blocks and each block has $H$ heads, we use the following metric to measure the attention sink of the whole LM: $\textrm{Sink}_{k}^{\epsilon}=\frac{1}{L}\sum_{l=1}^{L}\frac{1}{H}\sum_{h=1}^{H}\mathbb{I}(\alpha_{k}^{l\textrm{,}h}>\epsilon)$.
@@ 3.3 Attention sink under different inputs
Different data domains. We first explore the effects of input domains on attention sinks. The pile dataset (Gao et al., 2020), a regular dataset for LM pretraining, has 17 available data domains. As shown in Appendix B.2, input domains have negligible effects on our attention sink metric $\textrm{Sink}_{1}^{\epsilon}$.
Beyond natural languages. We also consider two ideal scenarios: (i) randomly sample $T$ tokens from the tokenizer vocabulary $\mathbb{V}$ to construct a sequence and (ii) randomly sample 1 token from the tokenizer $\mathbb{V}$ and repeat it $T$ times. We exclude the BOS token. As present in Table 1(Left), compared to natural language, random tokens have insignificant effects on the existence of attention sink. However, with repeated tokens, attention sink in Mistral (Jiang et al., 2023) and LLaMA models disappears. In Appendix B.1, we prove that for LMs with NoPE/relative PE/ALiBI/Rotary, if the first $T$ tokens are the same, their corresponding hidden states are the same. They all have massive activations, thus dispersing the attention sink. We also provide the closed form/upper bound for attention scores in these LMs through Propositions 1-4.
@@
Figure 4: (Left) Attention sink emerges in small LMs. (Middle) Dynamics of train/valid loss and $\textrm{Sink}_{1}^{\epsilon}$ during LM pre-training under the default setup. Attention sink emerges after certain optimization steps. (Right) Training loss (solid lines)/attention sink (dashed lines) dynamics of LMs using different learning rates. We observe that with smaller learning rates, attention sink tends to emerge after more optimization steps and be less obvious.

### v2

@@ Abstract
Auto-regressive Language Models (LMs) assign significant attention to the first token, even if it is not semantically important, which is known as attention sink. This phenomenon has been widely adopted in applications such as streaming/long context generation, KV cache optimization, inference acceleration, model quantization, and others. Despite its widespread use, a deep understanding of attention sink in LMs is still lacking. In this work, we first demonstrate that attention sinks exist universally in auto-regressive LMs with various inputs, even in small models. Furthermore, attention sink is observed to emerge during the LM pre-training, motivating us to investigate how optimization, data distribution, loss function, and model architecture in LM pre-training influence its emergence. We highlight that attention sink emerges after effective optimization on sufficient training data. The sink position is highly correlated with the loss function and data distribution. Most importantly, we find that attention sink acts more like key biases, storing extra attention scores, which could be non-informative and not contribute to the value computation. We also observe that this phenomenon (at least partially) stems from tokens’ inner dependence on attention scores as a result of softmax normalization. After relaxing such dependence by replacing softmax attention with other attention operations, such as sigmoid attention without normalization, attention sinks do not emerge in LMs up to 1B parameters. The code is available at https://github.com/sail-sg/Attention-Sink.
@@ 1 Introduction
Xiao et al. (2023b) showed that Large Language models (LLMs) allocate significant attention to the initial tokens, irrespective of their semantic relevance. This interesting phenomenon is termed as attention sink and has widespread applications, including streaming/long context generation (Xiao et al., 2023b; Han et al., 2024; Yang et al., 2024), KV cache optimization (Ge et al., 2023; Wan et al., 2024; Wu & Tu, 2024), efficient inference (Zhang et al., 2024b; Chen et al., 2024), model quantization (Liu et al., 2024b; Huang et al., 2024), and others.
A seminal of works attempted to understand attention sink. Among them, Cancedda (2024) clarified that attention sink primarily appears only on the first token. They attributed the phenomenon to the large norm of hidden states of the first token. This is referred to as massive activations (very few activations exhibit extremely large values compared to others) in Sun et al. (2024). Besides, Sun et al. (2024); Yu et al. (2024) observed that attention sink may also appear in several word tokens carrying limited semantic information and having no fixed position. Despite the above research efforts, a deep understanding of attention sink is still absent. Therefore, we conduct a comprehensive study to investigate when attention sink emerges. We defer full discussions on related work to Appendix A.
Based on open-sourced auto-regressive LMs, we show that the first token acts as biases: the angles between the first key and queries of other tokens are typically small, leading to attention sink. Then we find that attention sink universally exists in auto-regressive LMs across different inputs, even in the small models or with random token sequences. Additionally, attention sink is observed to emerge during the LM pre-training before continual instruction tuning (Ouyang et al., 2022). This motivates us to focus on the LM pre-training, whose objective can be formulated as:
@@ 2 Preliminaries on LMs and attention sink
$\bm{A}^{l\textrm{,}h}=\textrm{Softmax}\big{(}{\bm{Q}^{l\textrm{,}h}{\bm{K}^{l\textrm{,}h}}^{\top}}/{\sqrt{d_{h}}}+\bm{M}\big{)}\textrm{,}\,\,\,\,\bm{O}^{l}=\textrm{Concat}_{h=1}^{H}\left(\bm{A}^{l\textrm{,}h}\bm{V}^{l\textrm{,}h}\right)\bm{W}_{O}^{l}\textrm{,}$ (4)
@@
Positional embedding. NoPE (Kazemnejad et al., 2024) considered no explicit positional embedding (PE) in LMs, where $\bm{P}=\bm{0}$. When using absolute PE (Vaswani et al., 2017), $\bm{P}$ is a periodic function of token positions. Devlin et al. (2019); Brown et al. (2020) adopted a learnable PE, which means $\bm{P}$ is a learnable embedding of token positions. The dot product between each query and key meets $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{\top}$ when using the above three PEs. While for relative PE (Raffel et al., 2020), ALiBi (Press et al., 2021), Rotary (Su et al., 2024), they have $\bm{P}=\bm{0}$. Instead, they modify the dot product $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle$. For relative PE and ALiBi, $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{\top}+g(i-j)$, where $g(\cdot)$ is pre-defined function of the distance between two token positions. For Rotary, $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,j-i}\bm{k}_{j}^{\top}$, where $\bm{R}_{\Theta\textrm{,}\,(\cdot)}$ is a pre-defined rotation matrix. We include detailed formulations of the above PEs in Appendix B.
@@
Attention sink. Xiao et al. (2023b) revealed that LLMs allocate significant attention scores to specific token positions, e.g. the first token (not necessary to be a BOS token), resulting in “vertical” attention patterns. To represent this, we have the attention scores $\bm{A}_{i\textrm{,}\,1}^{l\textrm{,}\,h}\gg\textrm{mean}(\bm{A}_{i\textrm{,}\,j\neq 1}^{l\textrm{,}\,h})$.
Figure 2: In LLaMA3-8B Base, (Top) the first token has significantly larger $\ell_{2}$-norm of hidden states, but much smaller $\ell_{2}$-norm of keys and values than the mean of other tokens; (Bottom) cosine similarity instead of norm product contributes to attention sink. We delay more visualizations to Appendix C.3.
@@ 3.1 The first token acts as biases
Uniqueness of the first token. It is noted that the calculation of hidden states for the first token has no involvement of self-attention $\bm{h}_{1}^{l}=\textrm{FFN}(\textrm{LN}(\bm{o}_{1}^{l}+\bm{h}_{1}^{l-1}))+\bm{o}_{1}^{l}+\bm{h}_{1}^{l-1}$, where $\bm{o}_{1}^{l}=\textrm{LN}(\bm{h}_{1}^{l-1})\begin{bmatrix}\bm{W}^{l\textrm{,}1}&\bm{W}^{l\textrm{,}2}&\cdots&\bm{W}^{l\textrm{,}H}\\ \end{bmatrix}\bm{W}_{O}^{l}$. Therefore, $\bm{h}_{1}^{l}$, and corresponding queries/keys/values $\bm{k}_{1}^{l\textrm{,}h}=\textrm{LN}(\bm{h}_{1}^{l-1})\bm{W}_{K}^{l\textrm{,}h}$, $\bm{q}_{1}^{l\textrm{,}h}=\textrm{LN}(\bm{h}_{1}^{l-1})\bm{W}_{Q}^{l\textrm{,}h}$, $\bm{v}_{1}^{l\textrm{,}h}=\textrm{LN}(\bm{h}_{1}^{l-1})\bm{W}_{V}^{l\textrm{,}h}$ could be considered as the MLP output of input word embedding $\bm{x}_{1}\bm{W}_{E}$. Using LLaMA3-8B Base (Dubey et al., 2024), we show that from certain transformer block, e.g., $l=2$, the $\ell_{2}$-norm of $\bm{h}_{1}^{l}$ is significantly larger than that of other tokens $\bm{h}_{t\neq 1}$ in Figure 2(Top). This reproduces massive activations in Cancedda (2024); Sun et al. (2024). Despite the large $\ell_{2}$-norm of hidden states, we observe that the $\ell_{2}$-norm of keys and values of the first token is significantly smaller than that of other tokens in the same figure, which was also observed in Devoto et al. (2024); Guo et al. (2024b).
@@
For the $t$-th query, we know that $\bm{q}_{t}^{l\textrm{,}\,h}{\bm{k}_{1}^{l\textrm{,}\,h}}^{\top}$ typically has much larger values than $\bm{q}_{t}^{l\textrm{,}\,h}{\bm{k}_{j\neq 1}^{l\textrm{,}\,h}}^{\top}$, as visualized in Figure 2(Bottom). We further show that due to the different manifold of $\bm{k}_{1}^{l\textrm{,}\,h}$, the angles between $\bm{k}_{1}^{l\textrm{,}\,h}$ and $\bm{q}_{t}^{l\textrm{,}\,h}$ play an important role. Considering $\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{j}^{l\textrm{,}h}}^{\top}=\|\bm{q}_{t}^{l\textrm{,}h}\|\cdot\|\bm{k}_{j}^{l\textrm{,}h}\|\cdot\textrm{cos}(\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h})$, we visualize the cosine similarity between keys and values, and the product of $\ell_{2}$-norm between keys and values in Figure 2(Bottom). Although $\|\bm{q}_{t}^{l\textrm{,}h}\|\cdot\|\bm{k}_{1}^{l\textrm{,}h}\|$ is comparatively small, $\textrm{cos}(\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{1}^{l\textrm{,}h})$ is significantly large, leading to attention sink. This explains why attention sink exists despite the small $\ell_{2}$-norm of keys of the first token. To conclude, the first token leverages its keys to act as biases, thus minimizing the angles between $\bm{k}_{1}^{l\textrm{,}\,h}$ and $\bm{q}_{t}^{l\textrm{,}\,h}$ and exhibiting attention sink.
@@ 3.2 Measuring attention sink
Threshold-based metrics. Xiao et al. (2023b) showcased the appearance of attention sink by visualizing attention logits/scores in different heads/blocks. This leads to the intractability of measuring attention sink quantitatively due to the large number of attention heads and blocks. Therefore, we first explore the metrics to measure the attention sink. Within each head, we compute the importance scores for the $k$-th token $\alpha_{k}^{l\textrm{,}h}=\frac{1}{T-k+1}\sum_{i=k}^{T}\bm{A}_{i\textrm{,}\,k}^{l\textrm{,}\,h}$. We mainly focus on the first token $\alpha_{1}^{l\textrm{,}h}$. It is noted that $\frac{1}{T}\leq\alpha_{1}^{l\textrm{,}\,h}\leq 1$ since $\bm{A}_{1\textrm{,}\,1}^{l\textrm{,}\,h}=1$ and $0\leq\bm{A}_{i\neq 1\textrm{,}\,1}^{l\textrm{,}\,h}\leq 1$. Then we adopt a threshold-based metric, we consider a head has attention sink in the first token if $\alpha_{1}^{l\textrm{,}h}>\epsilon$. Considering that the whole model has $L$ blocks and each block has $H$ heads, we use the following metric to measure the attention sink of the whole LM: $\textrm{Sink}_{k}^{\epsilon}=\frac{1}{L}\sum_{l=1}^{L}\frac{1}{H}\sum_{h=1}^{H}\mathbb{I}(\alpha_{k}^{l\textrm{,}h}>\epsilon)$.
@@ 3.3 Attention sink under different inputs
Different data domains. We first explore the effects of input domains on attention sinks. The pile dataset (Gao et al., 2020), a regular dataset for LM pretraining, has 17 available data domains. As shown in Appendix C.2, input domains have negligible effects on our attention sink metric $\textrm{Sink}_{1}^{\epsilon}$.
Beyond natural languages. We also consider two ideal scenarios: (i) randomly sample $T$ tokens from the tokenizer vocabulary $\mathbb{V}$ to construct a sequence and (ii) randomly sample 1 token from the tokenizer $\mathbb{V}$ and repeat it $T$ times. As present in Table 1(Left), attention sink still exists when the inputs are random tokens instead of natural language. However, with repeated tokens, attention sink in Mistral (Jiang et al., 2023) and LLaMA models disappears. In Appendix C.1, we prove that for LMs with NoPE/relative PE/ALiBI/Rotary, if the first $T$ tokens are the same, their corresponding hidden states are the same. They all have massive activations, thus dispersing the attention sink. We also provide the closed form/upper bound for attention scores in these LMs through Propositions 1-4.
@@
Figure 4: (Left) Attention sink also emerges in small LMs. (Middle) Dynamics of train/valid loss and $\textrm{Sink}_{1}^{\epsilon}$ during LM pre-training under the default setup. Attention sink emerges after certain optimization steps. (Right) Training loss (solid lines)/attention sink (dashed lines) dynamics of LMs using different learning rates. We observe that with smaller learning rates, attention sink tends to emerge after more optimization steps and be less obvious.